OpenStudy (anonymous):

Can anyone give me a clear definition about H.C.F and L.C.M of polynomial? eg1) Find the H.C.F of \((x-1)(x ^{2}-4x+4)\) and \((x ^{2}-2x)(x ^{2}-1)\) . eg2) Find the L.C.M. of \(x ^{2}-1\) and \((x+1)^{2}\) .

OpenStudy (fibonaccichick666):

I'm assuming that is like highest common factor or something?

OpenStudy (anonymous):

how about LCM? lowest common mutiply?

OpenStudy (fibonaccichick666):

I have to assume it is just like finding the LCM of 3 and 4

OpenStudy (anonymous):

that's 12, isnt it?

OpenStudy (fibonaccichick666):

yea,

OpenStudy (anonymous):

find factors of x^2-4x + 4 and factors of x^2 -2x and x^2 -1 and then see what is H.C.F.

OpenStudy (anonymous):

(x-1)(x^2-4x+4)=(x-1)(x-2)^2

OpenStudy (anonymous):

(x^2-2x)(x^2-1)=x(x-2)(x-1)(x+1)

OpenStudy (anonymous):

then the HCF is (x-1)(x-2)???

OpenStudy (anonymous):

factors of x^2 -4x + 4 = (x-2)(x-2)

OpenStudy (fibonaccichick666):

you can have it BPD, I've never heard these terms before in my life, I'm using guesswork based off of other situations. I would say the highest common factor is the biggest quantity that divides (x-1)(x-2)^2

OpenStudy (anonymous):

factors of x^2 - 2x are (x)(x-2)

OpenStudy (fibonaccichick666):

so I agree with your answer kryton

OpenStudy (anonymous):

factors of x^2 - 1 (difference of two squares) are (x-1)(x+1)

OpenStudy (anonymous):

Now if you list your factors (no repeat): (x)(x-2)(x-1)(x-1) (x-1)(x-2)(x-2) the highest common factor in both of these is (x-2)

OpenStudy (fibonaccichick666):

I disagree there

OpenStudy (anonymous):

another way of looking at H.C.F. is something that will divide into both of these

OpenStudy (fibonaccichick666):

Now that I know for certain it is just like with integers, I would say the gcm is (x-2)(x-1)

OpenStudy (anonymous):

Explain please @FibonacciChick666

OpenStudy (fibonaccichick666):

as that is a factor of both

OpenStudy (fibonaccichick666):

can you divide the first factorization you gave evenly by (x-1)(x-2)?

OpenStudy (anonymous):

yes, you are right @FibonacciChick666, I missed that! yes it's (x-1)(x-2)

OpenStudy (fibonaccichick666):

\[\frac{(x)\not {(x-2)(x-1)}(x-1)}{\not{(x-1)(x-2)}}\]

OpenStudy (fibonaccichick666):

oh those were supposed to be slashes... but anyways, yea. I have just never seen HCM. In number theory even we used GCD

OpenStudy (fibonaccichick666):

greatest common divisor

OpenStudy (anonymous):

i agree with @FibonacciChick666

OpenStudy (anonymous):

thank you @FibonacciChick666 :) i have cleared the concept!

OpenStudy (fibonaccichick666):

so @kryton1212, the definition I will give you is that the HCM as you call it is the same as the greatest common divisor, which as the name implies, is the absolute biggest number you can divide a pair of polynomials by and have no remainder.

OpenStudy (fibonaccichick666):

np

OpenStudy (fibonaccichick666):

it was a group effort, I'm guessing you have to deal with common core?

OpenStudy (anonymous):

yup

OpenStudy (fibonaccichick666):

I am so sorry. I have to teach the idiocy of it in order for the kids I tutor to pass their class....

OpenStudy (anonymous):

i am always confused with polynomials

OpenStudy (anonymous):

oh never mind haha

OpenStudy (fibonaccichick666):

well, they really aren't that bad. There are a lot of things worse. What confuses you most?